OpenStudy (anonymous):

sketch each triangle and then solve the triangle using the law of sines. angle a is =30 degrees, angle c= 65 degrees and b=10

OpenStudy (anonymous):

First, you must find angle b. To find angle b, we simply subtract the other measures of the triangle from 180 (the sum of the degrees in a triangle.) Therefore, angle b= 85. Next, we will use the law of sines to find the remaining sides of the triangle. sin(A)/a=sin(B)/b. Therefore, (sin30)/a=(sin85)/10. Thus, a=5.0191 we can also solve for c in a similar fashion. sin(C)/c=sin(B)/b Therefore, (sin65)/c=(sin85)/10 so c=9.0977 So we have B=85 degrees, c=9.0977, and a=5.0191. Done :)

OpenStudy (anonymous):

wow thanks!!!